OpenStudy (shelby1290):

a) Does the point (2,-1) lie on the line y=-3x+5? b) Does the point (-4,10) lie on the parabola y=-2x^2 - 5x + 22?

jimthompson5910 (jim_thompson5910):

`a) Does the point (2,-1) lie on the line y=-3x+5?` the point (2,-1) means x = 2 and y = -1 replace x with 2 and y with -1. Then simplify each side y=-3x+5 y=-3*2+5 ... replacing x with 2 -1=-3*2+5 ... replacing y with -1 -1=-6+5 -1=-1 Since we got -1=-1, which is true, it makes the original equation true. So the point (2,-1) is definitely on the line y=-3x+5

jimthompson5910 (jim_thompson5910):

for part b), the same idea applies